OpenStudy (lucaz):

complex numbers, znot=conjugate z(znot)+z+znot=0

OpenStudy (lucaz):

I found a^2+2a+b^2=0

OpenStudy (freckles):

maybe we can use this \[z=a+bi =r e^{i \theta} \\ z'=a-bi=r e^{-i \theta}\]

OpenStudy (lucaz):

they are asking for the graph of z's that satisfy the equation

OpenStudy (freckles):

doesn't matter what r is any r will work that equation you can either try to solve as a quadratic or you can try to use \[\cos(\theta)=\frac{1}{2}(e^{i \theta} +e^{- i \theta})\]

OpenStudy (lucaz):

well, I haven't seen this expression yet. there is no other way to solve it?

OpenStudy (freckles):

wait which expression?

OpenStudy (lucaz):

e^(i*theta)

OpenStudy (freckles):

oh so no euler formula at all ?

OpenStudy (lucaz):

nope

OpenStudy (lucaz):

I started with z=a+bi

OpenStudy (lucaz):

and expended the equation

OpenStudy (freckles):

what ab out cos and sin?

OpenStudy (lucaz):

I didn't try the trigonometric form yet

OpenStudy (freckles):

\[z=r \cos(\theta)+i r \sin(\theta) \\ z'=r \cos(\theta)-i r \sin(\theta)\] can we use this ?

OpenStudy (freckles):

well you will see something cool happen if you do

OpenStudy (freckles):

and what I mean is you will get to use pythagorean identity

OpenStudy (lucaz):

ok, I will do it. thanks

OpenStudy (freckles):

you closed it does that mean you got it?

OpenStudy (lucaz):

yeah =]

OpenStudy (freckles):

cool stuff